Given that the volume of a cylinder is 150, and the radius of the cylinder is twice the height, find the surface Area of the cylinder.
I got 2766.75 as the answer but that is not correct

Answer by stanbon(75887)

Given that the volume of a cylinder is 150, and the radius of the cylinder is twice the height, find the surface Area of the cylinder.
----------------------
V = (pi)r^2
150 = (pi)r^2
r = 6.909..
---------------
Assuming the surface area does not include the area of the two bases
you get:
S = 2(pi)rh
S =2(pi)6.909*6.909/2
S = 6.909^2(pi)
S = 150 sq units
=================
Cheers,
Stan H.